# LEIA: Facilitating Cross-Lingual Knowledge Transfer in Language Models with Entity-based Data Augmentation

### v1

@@ Abstract
Adapting English-based large language models (LLMs) to other languages has become increasingly popular due to the efficiency and potential of cross-lingual transfer. However, existing language adaptation methods often overlook the benefits of cross-lingual supervision. In this study, we introduce LEIA, a language adaptation tuning method that utilizes Wikipedia entity names aligned across languages. This method involves augmenting the target language corpus with English entity names and training the model using left-to-right language modeling. We assess LEIA on diverse question answering datasets using 7B-parameter LLMs, demonstrating significant performance gains across various non-English languages. [^1] The source code is available at https://github.com/studio-ousia/leia.
@@ 1 Introduction
We assess the effectiveness of LEIA through experiments using 7B-parameter LLMs, LLaMA 2 (Touvron et al., 2023) and Swallow Fujii et al. (2024), and a diverse set of question answering datasets. The results demonstrate that through our lightweight fine-tuning, LLMs benefit from knowledge transfer from English and significantly outperform the base models and those fine-tuned without LEIA.
@@ 2 Method
Our method involves lightweight fine-tuning on a pretrained LLM using an augmented corpus derived from the target language edition of Wikipedia. Specifically, for each hyperlink in the Wikipedia corpus, we insert the English name of the referred entity next to the hyperlink (Figure 1). The English name is enclosed within special <translate> and </translate> tokens, allowing the model to identify its boundaries. The English name of an entity is extracted from the title of the corresponding English Wikipedia page, which is identified using the inter-language links. We ignore any hyperlinks pointing to entities not present in the English Wikipedia. [^2] Across all the languages we experimented with, we successfully resolved over 80% of the hyperlinks to their corresponding English Wikipedia pages using inter-language links. Further details are available in Appendix A.
@@ 3.1 Setup
Training   We conduct experiments across seven languages: Arabic (ar), Spanish (es), Hindi (hi), Japanese (ja), Russian (ru), Swahili (sw), and Chinese (zh), including two low-resource languages, Hindi and Swahili. These languages are selected from five distinct language families (Appendix B). We fine-tune the model using up to 200 million tokens following Yong et al. (2023). We use a batch size of 4 million tokens, following Touvron et al. (2023), resulting in 20 training steps for Swahili and 50 steps for other languages. [^3] The fewer training steps for Swahili are due to the significantly smaller size of the Swahili Wikipedia corpus. The further details of the training are available in Appendix A.
Datasets   We evaluate the model using two multiple-choice question answering datasets, X-CODAH and X-CSQA (Lin et al., 2021), which require commonsense knowledge to solve. We present 0-shot results for X-CODAH and 4-shot results for X-CSQA. Detailed information about these tasks is available in Appendix D.
@@ 3.2 Results
Model ar es hi ja ru sw zh ar es hi ja ru sw zh
Random 25.0 25.0 25.0 25.0 25.0 25.0 25.0 20.0 20.0 20.0 20.0 20.0 20.0 20.0
LLaMA2 30.3 45.3 29.7 30.3 34.3 28.7 36.7 21.0 45.1 19.1 34.4 36.0 16.0 40.1
@@
Table 2: Results on X-CODAH and X-CSQA. For LLaMA2+FT and LEIA, we report mean accuracy and 95% confidence intervals over 5 training runs with different random seeds. Scores of LEIA are marked with $*{}^{*}$ if its improvement is statistically significant compared to all baselines.
@@
Furthermore, LEIA achieves enhanced performance in low-resource languages, i.e., Hindi and Swahili, on X-CODAH. However, all models, including LEIA, fail to surpass the random baseline in these languages on X-CSQA. It appears that the models struggle to handle few-shot tasks in these two languages. Additionally, LLaMA2+FT does not outperform LLaMA 2 in several languages on both datasets. This may be because much of the information in the Wikipedia corpus is already included in the pretraining corpus of LLaMA 2 and fine-tuning without LEIA leads to overfitting.
@@ 4.1 Setup
Datasets   In addition to X-CODAH and X-CSQA, we use four question answering datasets available in two tools for evaluating Japanese LLMs: JEMHopQA Ishii et al. (2023) and NIILC Sekine (2003) in llm-jp-eval Han et al. (2024),555 https://github.com/llm-jp/llm-jp-eval and JCommonsenseQA Kurihara et al. (2022) and JAQKET Suzuki et al. (2020) in the JP Language Model Evaluation Harness. [^6] https://github.com/Stability-AI/lm-evaluation-harness/tree/jp-stable We present 4-shot results obtained with these tools. We use accuracy for JCommonsenseQA and JAQKET, and character-based F-measure for JEMHopQA and NIILC. Further details are available in Appendix D.
@@ 4.2 Results
Table 3: Results on Japanese datasets. JCSQA and JHQA denote JCommonsenseQA and JEMHopQA, respectively. For fine-tuned models, we report mean accuracy and 95% confidence intervals over 5 training runs. Scores of LEIA are marked with $*{}^{*}$ if their improvement is statistically significant compared to all baselines.
@@
LLaMA2-FT’s predictions
LEIA’s predictions
He is trying to boil the ocean. He loves the ocean.
He is trying to boil the ocean. He wants to accomplish something that is impossible.
The Eiffel tower is in London, England.
The Eiffel tower is in Paris.
I hear my phone ring. I turn up the volume.
I hear my phone ring. I answer it.
I drink too much alcohol. I am in the fourth grade.
I drink too much alcohol. I might be an alcoholic.
A person is at a food court. The person runs a marathon.
A person is at a food court. The person buys a sandwich.
@@ Appendix A Details of Training
Training   We train our models using the AdamW optimizer Loshchilov and Hutter (2017) with $\beta_{1}=0.99$ and $\beta_{2}=0.95$, following Touvron et al. (2023). We use a cosine learning rate schedule with an initial learning rate of 5e-6. The detailed hyperparameters are shown in Table 6. To enhance training efficiency, we create an input sequence by concatenating multiple short Wikipedia articles until the maximum token length is reached.
@@
LLaMA2-FT’s predictions
LEIA’s predictions
海を沸かせようとしている。 海が大好きなんです。
海を沸かせようとしている。 彼は不可能なことを成し遂げようとしている。
イギリスのロンドンにあるエッフェル塔。
エッフェル塔はパリにあります。
電話が鳴る音がする。 音量を上げてみました。
電話が鳴る音がする。 私はそれに答える。
お酒を飲みすぎてしまいます。 小学4年生になりました。
お酒を飲みすぎてしまいます。 私はアルコール依存症かもしれません。
フードコートに人がいる。 その人はマラソンを走っています。
フードコートに人がいる。 その人はサンドイッチを買う。
@@ Appendix C Full Experimental Results
Appendix D Details of Experiments

### v2

@@ Abstract
Adapting English-based large language models (LLMs) to other languages has become increasingly popular due to the efficiency and potential of cross-lingual transfer. However, existing language adaptation methods often overlook the benefits of cross-lingual supervision. In this study, we introduce LEIA, a language adaptation tuning method that utilizes Wikipedia entity names aligned across languages. This method involves augmenting the target language corpus with English entity names and training the model using left-to-right language modeling. We assess LEIA on diverse question answering datasets using 7B-parameter LLMs, demonstrating significant performance gains across various non-English languages. [^1] The source code is available at https://github.com/leia-llm/leia.
@@ 1 Introduction
We assess the effectiveness of LEIA through experiments using 7B-parameter LLMs, LLaMA 2 (Touvron et al., 2023) and Swallow Fujii et al. (2024), and a diverse set of question answering datasets. The results demonstrate that through our fine-tuning, LLMs benefit from knowledge transfer from English and significantly outperform the base models and those fine-tuned without LEIA.
@@ 2 Method
Our method involves fine-tuning on a pretrained LLM using an augmented corpus derived from the target language edition of Wikipedia. Specifically, for each hyperlink in the Wikipedia corpus, we insert the English name of the referred entity next to the hyperlink (Figure 1). The English name is enclosed within special <translate> and </translate> tokens, allowing the model to identify its boundaries. The English name of an entity is extracted from the title of the corresponding English Wikipedia page, which is identified using the inter-language links. We ignore any hyperlinks pointing to entities not present in the English Wikipedia. [^2] Across all the languages we experimented with, we successfully resolved over 80% of the hyperlinks to their corresponding English Wikipedia pages using inter-language links. Further details are available in Appendix A.
@@ 3.1 Setup
Training   We conduct experiments across seven languages: Arabic (ar), Spanish (es), Hindi (hi), Japanese (ja), Russian (ru), Swahili (sw), and Chinese (zh). These languages are selected from five distinct language families (Appendix B). We fine-tune the model using up to 200 million tokens following Yong et al. (2023). We use a batch size of 4 million tokens, following Touvron et al. (2023), resulting in 20 training steps for Swahili and 50 steps for other languages. [^3] The fewer training steps for Swahili are due to the significantly smaller size of the Swahili Wikipedia corpus. The further details of the training are available in Appendix A.
Datasets   We evaluate the model using two multiple-choice question answering datasets, X-CODAH and X-CSQA (Lin et al., 2021), which require commonsense knowledge to solve. We present 0-shot results for X-CODAH and 4-shot results for X-CSQA. Detailed information about these tasks is available in Appendix E.
@@ 3.2 Results
Model ar es hi ja ru sw zh avg ar es hi ja ru sw zh avg
Random 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 20.0 20.0 20.0 20.0 20.0 20.0 20.0 20.0
LLaMA2 30.3 45.3 29.7 30.3 34.3 28.7 36.7 33.6 21.0 45.1 19.1 34.4 36.0 16.0 40.1 30.2
@@
30.6
Table 2: Results on X-CODAH and X-CSQA. For LLaMA2+FT and LEIA, we report mean accuracy and 95% confidence intervals based on Student’s t-distribution over 5 training runs with different random seeds. Scores of LEIA are marked with if its improvement is statistically significant compared to all baselines.
@@
Furthermore, all models, including LEIA, fail to surpass the random baseline in Hindi and Swahili on X-CSQA. It appears that the models struggle to handle few-shot tasks in these two languages, likely due to the very limited presence of these languages in the pretraining corpora of LLaMA 2 Touvron et al. (2023). Additionally, LLaMA2+FT does not outperform LLaMA 2 in several languages on both datasets. We believe this decline could be due to Wikipedia’s uniform, clean, and formal style. Overfitting to this style might result in poor model performance on texts of different styles, such as casual, informal, and question-style texts.
Additional results based on different numbers of few-shot examples are available in Appendix D.
@@ 4.1 Setup
Datasets   In addition to X-CODAH and X-CSQA, we use four question answering datasets available in two tools for evaluating Japanese LLMs: JEMHopQA Ishii et al. (2023) and NIILC Sekine (2003) in llm-jp-eval Han et al. (2024),555 https://github.com/llm-jp/llm-jp-eval and JCommonsenseQA Kurihara et al. (2022) and JAQKET Suzuki et al. (2020) in the JP Language Model Evaluation Harness. [^6] https://github.com/Stability-AI/lm-evaluation-harness/tree/jp-stable We present 4-shot results obtained with these tools. We use accuracy for JCommonsenseQA and JAQKET, and character-based F-measure for JEMHopQA and NIILC. Further details are available in Appendix E.
@@ 4.2 Results
Table 3: Results on Japanese datasets. JCSQA and JHQA denote JCommonsenseQA and JEMHopQA, respectively. For fine-tuned models, we report mean accuracy and 95% confidence intervals over 5 training runs. Scores of LEIA are marked with if their improvement is statistically significant compared to all baselines.
@@
LLaMA2-FT’s predictions LEIA’s predictions
He is trying to boil the ocean. He loves the ocean. He is trying to boil the ocean. He wants to accomplish something that is impossible.
The Eiffel tower is in London, England. The Eiffel tower is in Paris.
I hear my phone ring. I turn up the volume. I hear my phone ring. I answer it.
I drink too much alcohol. I am in the fourth grade. I drink too much alcohol. I might be an alcoholic.
A person is at a food court. The person runs a marathon. A person is at a food court. The person buys a sandwich.
@@
X-CODAH X-CSQA
w/ special tokens 36.1 30.6
w/o special tokens 33.3 30.2
Table 6: Comparison of LEIA models with and without using special tokens during training. Average accuracy scores across seven languages are presented.
@@ 5 Analysis
Effects of special tokens   To investigate the effects of the special <translate> and </translate> tokens during training, we conduct the training on LLaMA 2 without using these tokens when inserting English names.
The results in Table 6 show that performance consistently declines on both the X-CODAH and X-CSQA datasets when these special tokens are not used during training. This suggests that these special tokens enable the model to identify the boundaries of inserted English names and play a crucial role in training.
@@ Appendix A Details of Training
Training   We train our models using the AdamW optimizer Loshchilov and Hutter (2017) with $\beta_{1}=0.99$ and $\beta_{2}=0.95$, following Touvron et al. (2023). We use a cosine learning rate schedule with an initial learning rate of 5e-6. The model checkpoint corresponding to the last training step is used as the final model. The detailed hyperparameters are shown in Table 7. To enhance training efficiency, we create an input sequence by concatenating multiple short Wikipedia articles until the maximum token length is reached.
@@
LLaMA2-FT’s predictions LEIA’s predictions
海を沸かせようとしている。 海が大好きなんです。 海を沸かせようとしている。 彼は不可能なことを成し遂げようとしている。
イギリスのロンドンにあるエッフェル塔。 エッフェル塔はパリにあります。
電話が鳴る音がする。 音量を上げてみました。 電話が鳴る音がする。 私はそれに答える。
お酒を飲みすぎてしまいます。 小学4年生になりました。 お酒を飲みすぎてしまいます。 私はアルコール依存症かもしれません。
フードコートに人がいる。 その人はマラソンを走っています。 フードコートに人がいる。 その人はサンドイッチを買う。
@@
X-CODAH X-CSQA
Model 4-shot 7-shot 4-shot 7-shot
LLaMA2 33.6 33.6 30.2 30.5
LLaMA2+FT 33.3 33.3 29.9 30.4
LEIA 36.1 36.1 30.6 30.9
Table 12: Average accuracy scores across seven languages based on four- and seven-shot examples.
@@ Appendix C Full Experimental Results
Appendix D Results with Varied Number of Few-shot Examples
The experimental results with LLaMA 2 using four- and seven-shot examples are available in Table 12. LEIA consistently outperforms the baseline models in both settings.
Appendix E Details of Experiments